OpenStudy (anonymous):

solve the system of linear equations 2x+3y=2 x+6y=4

OpenStudy (blurbendy):

what does x equal on the bottom equation

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

firstm ultiply the bottom equation to get the x's cancle out with -2 so you got 2x+3y=2 -2x-12y=-8 the x's cancel and you get -9y=-6 divide by -9 and you get \[y=\frac{ 2 }{ 3 }\] do same for y or just plug in 2/3 into one the equations to get y.

OpenStudy (anonymous):

i mean to get x not y

OpenStudy (anonymous):

how did you get -2

OpenStudy (anonymous):

you mean how did i come up with multiplying by -2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well look at both lines 2x+3y=2 x+6y=4 you want one the variables to cancel out by multiplying by an umber on one or both lines. The easiest way is to multiply the 2nd by -2 so you have 2x in first line and -2x in 2nd line. They both cancel each other and you got only a y variable and what it equals to xsolve for y.

OpenStudy (anonymous):

so by putting -2x it cancel 2x so -2 would not change the whole equation -2x+12y=8

OpenStudy (anonymous):

not exactly if your multiply a number into the line you multiply each term so it becomes -2x-12y=-8

OpenStudy (anonymous):

oh yea forgot about the negative

OpenStudy (anonymous):

yea but if you get -2x-12y=-8 and line the lines up again: 2x+3y=2 -2x-12y=-8 notice the 2x and -2x cancel and combine the rest you get -9y=-6 and solve for y.